Find the geometric mean between 4 sqareroot 3 and 10 sqareroot 3

Answer by Boreal(15235)

geometric mean of two things is the square root of the product of those two things
square root of 4sqrt(3)*10sqrt(3)=40*3=120
sqrt(120)=2 sqrt (30)